# Concerns and Values in Human-Robot Interactions: A Focus on Social Robotics

### v1

@@ B.4 Scenario 4
Sarah wakes up to the gentle sound of her favorite music playing softly through her bedroom speakers. She stretches and says, “Good morning, HomeAssist.” The system responds with a cheerful tone, “Good morning, Sarah! The weather today is sunny with a high of 75 degrees. You have a doctor’s appointment at 10 AM and a call with your sister at 3 PM. Would you like to review your schedule?”
She nods, and the system provides a detailed rundown of her day. Afterward, it offers, “Would you like to start your morning routine now?” Sarah agrees, and HomeAssist guides her to the bathroom, reminding her of the location of her toothbrush and other essentials.
Breakfast is Sarah’s next goal. She heads to the kitchen where the Smart Kitchen Assistant is ready to help. “What would you like to cook today, Sarah?” it asks. “How about some scrambled eggs and toast?” she replies. The assistant provides step-by-step voice instructions, alerting her when the pan is heated and when to flip the eggs. Sensors in the kitchen monitor the stove and other appliances, ensuring they are safely used. If the assistant detects any safety issue, like a burner left on, it promptly alerts Sarah and offers to turn it off automatically.
@@ Appendix C HRI Value Compass on paper
The latest paper version of the HRI Value Compass is available at https://doi.org/10.5281/zenodo.14002841.
@@
Appendix D Initial values from Phase 1
Appendix E Results from Phase 2: codes and themes
We first automatically transcribed the recordings of the focus group and we then manually checked them. Using the framework method for qualitative analysis [13] (i.e., using the values identified in our scoping review as main themes and the topics as main codes), we summarise the focus group findings as follows. In bold we highlight the values emerged, and in italic the corresponding topics,
E.1 Activity 1: scenario analysis
@@
Under the umbrella value of empowerment, data access emerged as a significant concern in scenarios 1 and 3. Specifically, P4 highlighted the importance of preserve data protection for children in scenario 3 and P1 pointed out that disclaiming upfront who is going to access the data collected in scenario 1 was of extremely importance. Analogously, manipulation has been identified as a source of concern for scenarios 2,3 and 4. P2 highlighted the complexity of scenario 1, which involves children, and argued that children should be meaningfully consulted about their values. P2 also questioned the appropriate “level of paternalism” to apply in such a complex scenario. P4 raised concerns about the potential influence of the robot on children’s behavior, particularly in terms of “nudging and manipulation”. P6 pointed out that in scenario 2, the child, Luke, might not have the capacity to “understand or consent to the use of the robot”, leaving the decision-making responsibility to the caregiver. She also emphasised that, particularly in scenario 3, it is crucial to establish clear guidelines on the extent to which students or learners can exercise agency in deciding the types and amounts of interactions they will engage in.
@@
Participants identified privacy as a key concern in scenarios 1 and 4. P1 questioned whether the Kim in scenario 1 is “truly able to consent to data collection and use,” given her limited “understanding of the technology.” P3 noted that privacy was the main issue in scenario 4, as it takes place in a “private environment where legislation is likely more relaxed”. Also in terms of intimacy, P1 suggested that there could be a “panopticon effect” in scenario 1 on the Kim’s behaviour. Assuming that Kim “is aware she is always being observed” and that various stakeholders could access her data, P1 questioned whether Kim might “self-police” her behavior, acting in ways she believes would be acceptable to potential observers. P1 also argued that the child could “effectively lose access to her private life” due to the robot’s constant presence and monitoring. The problem of profiling has emerged in all four scenarios. In scenario 1, P1 questioned whether the data collection is always in Kim’s “best interest.” For instance, P1 wondered if the data could be shared with Kim’s insurance company or sold to third parties. In scenario 2, P1 noted that these concerns are “even more pronounced” because Luke is a child. He argued that when forming a relationship with the robot, Luke may not be able to fully “consent to how his data is being used,” suggesting that such decisions should be left to his parents or guardians. In scenario 3, P5 introduced the concept of “datification” (i.e., reducing people to data) emphasising that designers and researchers should avoid collecting data on children’s abilities in ways that could lead to discrimination. Instead, they stressed the importance of preventing the datification of childhood altogether. In scenario 4, participants pointed out that a home environment may be more sensitive because it is not restricted to a medical setting where stricter rules may apply, increasing the potential risks and harms. They specifically mentioned that in a domestic context, data-capturing systems are particularly vulnerable to “misuse by domestic abusers.”
@@
Regarding welfare, P1 and P3 emphasized that in scenario 1, safety related to Kim’s physical well-being in health contexts is extremely important. P3 also suggested that there might be a “slight restriction” on Kim’s autonomy when the robot provides comfort to her during moments of distress. P5 also highlighted the importance of medical values, especially in scenario 1, for ensuring a beneficial interactions for the target user.
@@
Participants raised the issue of equality (referred to as equity), which could lead to both positive and negative outcomes. In scenario 1, P2 pointed out that while the robot observes Kim’s behavior and responds accordingly, it is unclear how the robot ensures that the values it displays truly reflect Kim’s desires and interests. For scenario 2, P1 noted that the robot will “inevitably make certain assumptions” leading to exclusion based on its programming about what constitutes “normal” behavior, and these assumptions could have problematic effects. This highlights that social behavior may vary depending on different contextual factors, such as culture. P1 expressed several concerns about potential algorithmic bias in EduBot in scenario 3. For example, he noted that the robot might struggle to accurately identify emotional states in children with different cognitive abilities or socioeconomic backgrounds, or it might “treat different kinds of students differently” in undesirable ways, such as reinforcing gender stereotypes.
@@
E.2 Activity 2: values revision
@@
P4 noted that autonomy and empowerment are closely related, suggesting that the concept of empowerment should be integrated into autonomy, which he defined as the “capacity of a person to make decisions independently and rationally.” P4 emphasized that while independence from others is important, it is not the same as autonomy, which involves making decisions without external impositions. He argued that manipulation should also be considered under autonomy because it impacts decision-making. P1 agreed that the value of autonomy needed clarification. He proposed distinguishing between two types of autonomy: rational decision-making and agency, which involves having options and the ability to act on them [72]. P5 pointed out that autonomy should be considered also in relation with humans. It is possible to undermine a person’s autonomy even when it is independent of other humans. In the scenario 4, for example, the individual does not rely on other humans for support or decision-making. P4, P1, and P5 agreed that honour should be redefined as dignity and should not include agency. P4 also suggested that manipulation should be categorised under autonomy rather than empowerment and recommended adding the concept of nudging, which was previously missing. Regarding privacy, P1 pointed out that having a robot present at all times could undermine autonomy and dignity, as constant surveillance might lead to persistent worry. He also clarified that fairness encompasses not just equal treatment but also addresses algorithmic bias. P5 emphasised that privacy considerations can have positive aspects when data is collected in compliance with established regulations, as such data can be valuable for improving technology. P3 emphasized the critical importance of the value and ethics of care, suggesting it should be explicitly recognized rather than being subsumed under welfare. He proposed rewording equality as equity. P1 felt that truth was not the best term to use, arguing that accountability is more relevant. He also recommended adding security to address concerns about potential control and misuse of the system and data. Additionally, P1 highlighted the importance of power, particularly in terms of who makes decisions, such as a single corporation controlling all robots. Lastly, P4 argued that kindness should be considered a virtue rather than a value and suggested removing it from the list. P5 emphasised the critical importance of contextualizing values, noting that they are situation-dependent and must be prioritized based on the specific circumstances. P6 commented on the overall categorization of values and topics, suggesting that some topics could themselves be considered values, as noted in the ethics of AI literature [42].
E.3 Activity 3: contextual information and relationship with values
@@
Participants believed that empowerment is strongly related to the “who” (specifically the role of the robot) and the “what” (the users’ goals). P1 suggested that this moral principle depends on whether the robot is replacing a human role, enhancing connections, or enabling new activities. P3 emphasized that to truly empower users, the robot’s primary task should not be to replace humans but to support them. For example, in math learning scenarios, the robot’s goal should focus on teaching skills rather than the subject itself. P5 raised the issue of situations where an elderly individual has written advanced directives, and the robot is responsible for calling for assistance. The question is how to handle these autonomous decisions made by the robot. P6 highlighted that the users are not necessarily able to fully understand what are their needs and what they want. She also pointed out that the users’ desires may differ from what the caregiver believes is best for them, which could create tension when defining values. Privacy was considered as closely related to the “where,” “how,” and “what contexts” in which the robot is used. P3 emphasized that using a robot at home is “completely different” from using it in a public space. P1 pointed out that how the robot is used is also crucial; for instance, if the robot is connected to other devices, it could impact the flow of data in and out of the robot. P4 highlighted the importance of the social and contextual environment in which the robot operates. For example, if a person “does not want others to know they are using the robot,” privacy should be maintained in all use cases. P5 suggested that the current legal scenario is important to take into account when complying with regulation frameworks. For welfare, participants believed it was linked to the “where,” “when,” and “why” the robot is used. P1 emphasized that safety must be ensured in all environments, particularly when robot navigation is involved. P4 highlighted that welfare is crucial when there are constraints on when the robot should be used; for example, ensuring that a robot is not used by a child at night. P4 also noted that the ultimate purpose of using a robot should be to promote welfare by making “life easier” and “people happier”. Participants discussed equality, also referred to as equity, as being related to the “who,” “where,” “how,” and “what constraints” when using the robot. P1 noted that equity is connected to demographics and that the environment (”where”) significantly impacts the robot’s performance. For example, if a robot overheats, its location might affect how easily it can be repaired. P3 added that the way a robot is used can limit access; for instance, if it requires another device to operate, some people may not be able to use it. Regarding constraints, P1 expressed concerns about manipulation, deception, and other “cognitive limitations” that could make people more susceptible to a robot’s behavior. P5 thought it was important to distinguish between different geographical regions and cultures to understand the values that predominate in a given context. For example, in the EU, there’s a strong focus on prioritizing the rights of individuals, which influences how technology, like autonomous robots, is used and developed. P6 highlighted that for the definition of “who” designers should wonder if they are you choosing the right kind of people by considering at-risk groups, vulnerability, and the inclusion or exclusion of stakeholders.
@@
Autonomy was considered to be related to the “who” and “what” of using a robot. P4 emphasized that when designing a robot, the focus should be on enhancing the user’s autonomy, not the robot’s. It is crucial to protect vulnerable populations, such as children and the elderly, and support their independence. P4 viewed the robot as a tool, an “enabler” that helps users achieve autonomy, and suggested implementing age-related restrictions on its use. P1 highlighted the importance of the robot assisting in decision-making that aligns with the person’s best interests and enabling them to do things they otherwise could not. Participants viewed honour as related to cultural dignity and social norms concerning how robots should behave with the users they interact with. Truth was discussed in relation to the “where” and “why” a robot is used. P1 noted that accountability and control over a robot can vary depending on the scenario. Similarly, P4 emphasized the importance of being transparent about the reasons for using the robot.

### v2

@@ B.4 Scenario 4
Sarah wakes up to the gentle sound of her favourite music playing softly through her bedroom speakers. She stretches and says, “Good morning, HomeAssist.” The system responds with a cheerful tone, “Good morning, Sarah! The weather today is sunny with a high of 75 degrees. You have a doctor’s appointment at 10 AM and a call with your sister at 3 PM. Would you like to review your schedule?”
She nods, and the system provides a detailed rundown of her day. Afterwards, it offers, “Would you like to start your morning routine now?” Sarah agrees, and HomeAssist guides her to the bathroom, reminding her of the location of her toothbrush and other essentials.
Breakfast is Sarah’s next goal. She heads to the kitchen where the Smart Kitchen Assistant is ready to help. “What would you like to cook today, Sarah?” it asks. “How about some scrambled eggs and toast?” she replies. The assistant provides step-by-step voice instructions, alerting her when the pan is heated and when to flip the eggs. Sensors in the kitchen monitor the stove and other appliances, ensuring they are used safely. If the assistant detects any safety issue, like a burner left on, it promptly alerts Sarah and offers to turn it off automatically.
@@ Appendix C HRI Value Compass on paper
The latest paper version of the HRI Value Compass is available at https://doi.org/10.5281/zenodo.14002841. Note that the font is small as these are supposed to be printed on A4 or A3 paper.
@@
Appendix D Results from Phase 2: codes and themes
We first automatically transcribed the recordings of the focus group and we then manually checked them. Using the framework method for qualitative analysis [braun2012thematic] (i.e., using the values identified in our scoping review as main themes and the topics as main codes), we summarise the focus group findings as follows. In bold we highlight the values that emerged, and in italic the corresponding topics,
D.1 Activity 1: scenario analysis
@@
Under the umbrella value of empowerment, data access emerged as a significant concern in scenarios 1 and 3. Specifically, P4 highlighted the importance of preserving data protection for children in scenario 3 and P1 pointed out that disclaiming upfront who is going to access the data collected in scenario 1 was of extreme importance. Analogously, manipulation has been identified as a source of concern for scenarios 2,3 and 4. P2 highlighted the complexity of scenario 1, which involves children, and argued that children should be meaningfully consulted about their values. P2 also questioned the appropriate “level of paternalism” to apply in such a complex scenario. P4 raised concerns about the potential influence of the robot on children’s behaviour, particularly in terms of “nudging and manipulation”. P6 pointed out that in scenario 2, the child, Luke, might not have the capacity to “understand or consent to the use of the robot”, leaving the decision-making responsibility to the caregiver. She also emphasised that, particularly in scenario 3, it is crucial to establish clear guidelines on the extent to which students or learners can exercise agency in deciding the types and amounts of interactions they will engage in.
@@
Participants identified privacy as a key concern in scenarios 1 and 4. P1 questioned whether the Kim in scenario 1 is “truly able to consent to data collection and use,” given her limited “understanding of the technology.” P3 noted that privacy was the main issue in scenario 4, as it takes place in a “private environment where legislation is likely more relaxed”. Also in terms of intimacy, P1 suggested that there could be a “panopticon effect” in scenario 1 on the Kim’s behaviour. Assuming that Kim “is aware she is always being observed” and that various stakeholders could access her data, P1 questioned whether Kim might “self-police” her behavior, acting in ways she believes would be acceptable to potential observers. P1 also argued that the child could “effectively lose access to her private life” due to the robot’s constant presence and monitoring. The problem of profiling has emerged in all four scenarios. In scenario 1, P1 questioned whether the data collection is always in Kim’s “best interest.” For instance, P1 wondered if the data could be shared with Kim’s insurance company or sold to third parties. In scenario 2, P1 noted that these concerns are “even more pronounced” because Luke is a child. He argued that when forming a relationship with the robot, Luke may not be able to fully “consent to how his data is being used,” suggesting that such decisions should be left to his parents or guardians. In scenario 3, P5 introduced the concept of “datification” (i.e., reducing people to data), emphasising that designers and researchers should avoid collecting data on children’s abilities in ways that could lead to discrimination. Instead, they stressed the importance of preventing the datification of childhood altogether. In scenario 4, participants pointed out that a home environment may be more sensitive because it is not restricted to a medical setting where stricter rules may apply, increasing the potential risks and harms. They specifically mentioned that in a domestic context, data-capturing systems are particularly vulnerable to “misuse by domestic abusers.”
@@
Regarding welfare, P1 and P3 emphasised that in scenario 1, safety related to Kim’s physical well-being in health contexts is extremely important. P3 also suggested that there might be a “slight restriction” on Kim’s autonomy when the robot provides comfort to her during moments of distress. P5 also highlighted the importance of medical values, especially in scenario 1, for ensuring beneficial interactions for the target user.
@@
Participants raised the issue of equality (referred to as equity), which could lead to both positive and negative outcomes. In scenario 1, P2 pointed out that while the robot observes Kim’s behaviour and responds accordingly, it is unclear how the robot ensures that the values it displays truly reflect Kim’s desires and interests. For scenario 2, P1 noted that the robot will “inevitably make certain assumptions” leading to exclusion based on its programming about what constitutes “normal” behaviour, and these assumptions could have problematic effects. This highlights that social behaviour may vary depending on different contextual factors, such as culture. P1 expressed several concerns about potential algorithmic bias in EduBot in scenario 3. For example, he noted that the robot might struggle to accurately identify emotional states in children with different cognitive abilities or socioeconomic backgrounds, or it might “treat different kinds of students differently” in undesirable ways, such as reinforcing gender stereotypes.
@@
D.2 Activity 2: values revision
@@
P4 noted that autonomy and empowerment are closely related, suggesting that the concept of empowerment should be integrated into autonomy, which he defined as the “capacity of a person to make decisions independently and rationally.” P4 emphasised that while independence from others is important, it is not the same as autonomy, which involves making decisions without external impositions. He argued that manipulation should also be considered under autonomy because it impacts decision-making. P1 agreed that the value of autonomy needed clarification. He proposed distinguishing between two types of autonomy: rational decision-making and agency, which involves having options and the ability to act on them [prunkl2022human]. P5 pointed out that autonomy should also be considered in relation to humans. It is possible to undermine a person’s autonomy even when it is independent of other humans. In scenario 4, for example, the individual does not rely on other humans for support or decision-making. P4, P1, and P5 agreed that honour should be redefined as dignity and should not include agency. P4 also suggested that manipulation should be categorised under autonomy rather than empowerment and recommended adding the concept of nudging, which was previously missing. Regarding privacy, P1 pointed out that having a robot present at all times could undermine autonomy and dignity, as constant surveillance might lead to persistent worry. He also clarified that fairness encompasses not just equal treatment but also addresses algorithmic bias. P5 emphasised that privacy considerations can have positive aspects when data is collected in compliance with established regulations, as such data can be valuable for improving technology. P3 emphasised the critical importance of the value and ethics of care, suggesting it should be explicitly recognised rather than being subsumed under welfare. He proposed rewording equality as equity. P1 felt that truth was not the best term to use, arguing that accountability is more relevant. He also recommended adding security to address concerns about potential control and misuse of the system and data. Additionally, P1 highlighted the importance of power, particularly in terms of who makes decisions, such as a single corporation controlling all robots. Lastly, P4 argued that kindness should be considered a virtue rather than a value and suggested removing it from the list. P5 emphasised the critical importance of contextualising values, noting that they are situation-dependent and must be prioritised based on the specific circumstances. P6 commented on the overall categorisation of values and topics, suggesting that some topics could themselves be considered values, as noted in the ethics of AI literature [hagendorff2020ethics].
D.3 Activity 3: contextual information and relationship with values
@@
Participants believed that empowerment is strongly related to the “who” (specifically the role of the robot) and the “what” (the users’ goals). P1 suggested that this moral principle depends on whether the robot is replacing a human role, enhancing connections, or enabling new activities. P3 emphasised that to truly empower users, the robot’s primary task should not be to replace humans but to support them. For example, in math learning scenarios, the robot’s goal should focus on teaching skills rather than the subject itself. P5 raised the issue of situations where an elderly individual has written advanced directives, and the robot is responsible for calling for assistance. The question is how to handle these autonomous decisions made by the robot. P6 highlighted that the users are not necessarily able to fully understand what their needs are and what they want. She also pointed out that the users’ desires may differ from what the caregiver believes is best for them, which could create tension when defining values. Privacy was considered as closely related to the “where,” “how,” and “what contexts” in which the robot is used. P3 emphasised that using a robot at home is “completely different” from using it in a public space. P1 pointed out that how the robot is used is also crucial; for instance, if the robot is connected to other devices, it could impact the flow of data in and out of the robot. P4 highlighted the importance of the social and contextual environment in which the robot operates. For example, if a person “does not want others to know they are using the robot,” privacy should be maintained in all use cases. P5 suggested that the current legal scenario is important to take into account when complying with regulatory frameworks. For welfare, participants believed it was linked to the “where,” “when,” and “why” the robot is used. P1 emphasised that safety must be ensured in all environments, particularly when robot navigation is involved. P4 highlighted that welfare is crucial when there are constraints on when the robot should be used; for example, ensuring that a robot is not used by a child at night. P4 also noted that the ultimate purpose of using a robot should be to promote welfare by making “life easier” and “people happier”. Participants discussed equality, also referred to as equity, as being related to the “who,” “where,” “how,” and “what constraints” when using the robot. P1 noted that equity is connected to demographics and that the environment (”where”) significantly impacts the robot’s performance. For example, if a robot overheats, its location might affect how easily it can be repaired. P3 added that the way a robot is used can limit access; for instance, if it requires another device to operate, some people may not be able to use it. Regarding constraints, P1 expressed concerns about manipulation, deception, and other “cognitive limitations” that could make people more susceptible to a robot’s behaviour. P5 thought it was important to distinguish between different geographical regions and cultures to understand the values that predominate in a given context. For example, in the EU, there’s a strong focus on prioritising the rights of individuals, which influences how technology, like autonomous robots, is used and developed. P6 highlighted that for the definition of “who” designers should wonder if they are choosing the right kind of people by considering at-risk groups, vulnerability, and the inclusion or exclusion of stakeholders.
@@
Autonomy was considered to be related to the “who” and “what” of using a robot. P4 emphasised that when designing a robot, the focus should be on enhancing the user’s autonomy, not the robot’s. It is crucial to protect vulnerable populations, such as children and the elderly, and support their independence. P4 viewed the robot as a tool, an “enabler” that helps users achieve autonomy, and suggested implementing age-related restrictions on its use. P1 highlighted the importance of the robot assisting in decision-making that aligns with the person’s best interests and enabling them to do things they otherwise could not. Participants viewed honour as related to cultural dignity and social norms concerning how robots should behave with the users they interact with. Truth was discussed in relation to the “where” and “why” a robot is used. P1 noted that accountability and control over a robot can vary depending on the scenario. Similarly, P4 emphasised the importance of being transparent about the reasons for using the robot.